What will be the composition of two functions f ° g from R to R,

(a)

f(x) = 2x + 3, g(x) = 3x + 2

(b)

f(x) = 2x + 3, g(x) = sin (x)

(c)

f(x) = sin (x), g(x) = 2x + 3

(d)

f(x) = sin (x), g(x) = x2

(e)

f(x) = |x|, g(x) = 2x + 3


1
Expert's answer
2021-07-19T16:57:51-0400

a)


"f\circ g=2(3x+2)+3"

"f\circ g=6x+7"

b)


"f\circ g=2\sin x+3"

c)


"f\circ g=\sin (2x+3)"

d)


"f\circ g=\sin (x^2)"

e)


"f\circ g=|2x+3|"